OpenStudy (anonymous):

Factor each trinomial below. Please show your work and check your answer. x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20 a2 + 12ab + 27b2 2a2 + 30a + 100

OpenStudy (anonymous):

Check your answers against the following list:\[\left\{(x-5) (x-3),(a-5) (a+4),a^2-5 a-20,(a+3 b) (a+9 b),2 (a+5) (a+10)\right\} \]The third expression cannot be factored.

OpenStudy (jiteshmeghwal9):

x^2-8x+15 =>x^2+3x-5x+15 =>x(x+3)-5(x-3) =>(x-5)(x^2-9)

OpenStudy (anonymous):

a2 – a – 20 a2-5a+4a-20 (a+4)(a-5) for a2 – 5a – 20 use (-b+b2-4ac)/2a ,(-b-b2-4ac)/2a for a2 + 12ab + 27b2=(a+3b)(a+9b) for 2a2 + 30a + 100=(2a+10)(a+10)